# Standalone 16-bit Neural Network Training: Missing Study for Hardware-Limited Deep Learning Practitioners

### v5

@@ Abstract
Comprehensive data and detailed instructions for reproducing our results are publicly available at: https://github.com/YUNBLAK/standalone_16bits_nn.
Keywords: Deep Learning, Mixed Precision, Floating-Point, Hardware-Limited
@@ 1 Introduction
As machine learning models grow in complexity and size, the demand for computational resources such as memory and processing power has become a critical challenge. To address this, the mixed precision technique has been developed to balance resource consumption and computational efficiency [38, 53, 7, 55]. This approach strategically employs different numerical precision types during forward and backward propagation, weight storage, or inference. For instance, GPT-3 was trained with mixed precision of 32-bit and 16-bit formats [2]; Meta’s LLaMA 3 employs FP8 precision in both the inference and forward pass training phases [14, 44]. Additionally, much work involves quantizing pre-trained, full-precision models to lower precisions, such as int8 or even 1-bit, particularly for weight storage and inference [16, 56].
The trend toward mixed precision is closely linked to the development of lower-precision numerical formats. The ongoing ML arithmetic standard IEEE P3109, for example, has formalized seven 8-bit floating-point formats [1]; The Nvidia Blackwell chip is pioneering 4-bit precision; and research is even exploring the feasibility of ultra-low-bit neural networks for LLMs [37, 35]. Since these lower precision formats typically cannot function effectively on their own, mixed precision has become a must, now a dominant practice in machine learning.
Paradoxically, while mixed precision is designed to mitigate resource pressures in advanced computing, access to this technology often remains elusive for many machine learning practitioners facing hardware limitations. Table 1 shows some major GPUs in Nvidia and their hardware support for floating-point. Both FP32 and FP16 are supported across all GPUs listed. Support for bfloat16 begins with Nvidia’s Ampere architecture, whereas FP8 is supported starting from the Hopper architecture, and FP4 is available from the most recent Blackwell architecture [41, 42]. Note that, high-end, costly equipment such as Nvidia’s Hopper GPUs or BrainChip’s neuromorphic chips is not readily available to all.
One might argue that practitioners can have access to lower precision through software simulation and that some non-floating-point formats such as int8 or fixed-point do not require hardware acceleration [48, 9]. These options can be beneficial if used in weight storage or inference, but they are not practical due to the significant slowdowns in training time when relying on software simulation or without hardware acceleration. Therefore, for those practitioners constrained by hardware, the options for training models (to be differentiated from inference and storage) are typically limited to using 32-bit, 16-bit, or a combination of the two in mixed precision.
Among the three precisions available to hardware-constrained researchers, 16-bit or its mixed precision with 32-bit should be preferred theoretically, as they offer memory and performance advantages while being well-supported by modern hardware accelerators and frameworks. Yet, surprisingly, there’s not much information in current studies about the ability of the 16-bit formats to work by themselves. Despite the prevailing belief among machine learning practitioners regarding the benefits of 16-bit training, there has not been a definitive study specifically addressing its efficacy as a standalone solution. As is well known, there are two types of 16-bit precision: the IEEE binary16 and Google Brain’s bfloat16. Despite the growing interest in bfloat16, this work focuses on IEEE binary16 due to its wider accessibility, particularly for hardware-constrained practitioners. This decision is based on the fact that IEEE binary16 is more commonly available than bfloat16. Henceforth, we will refer to the 16-bit precision standardized by IEEE as 16-bit or FP16 interchangeably.
Table 1: Supported floating-point precisions across major NVIDIA GPU architectures. [41, 42]
@@
The missing study in standalone functionality of 16bit is likely due to two factors. Firstly, those not limited by hardware constraints are progressively shifting towards lower precision models such as 8-bit or 4-bit. Secondly, researchers experimenting with pure 16-bit neural networks observed numerical instability, which led to the proposal of mixed-precision techniques to address this issue [53]. So, 16-bit precision may have gained a reputation for being insufficient as a standalone solution. Indeed, direct attempts to train neural networks using 16-bit precision in frameworks such as TensorFlow or PyTorch typically fail. However, as will be discussed later, these issues are likely due to incorrect parameter settings. In particular, the default optimizer is Tensorflow and Pytorch is Adam, which updates weights based on a formula that relies a certain $\epsilon$ in the denominator (see Sect. 5.3 for details). In TensorFlow, the default value of $\epsilon$ is set at 1E-7, which shows to be too small for FP16 training. Meanwhile, PyTorch sets the default $\epsilon$ value at 1E-8, which effectively rounds to zero in FP16, as its smallest representable positive number is approximately 5.96E-8. Thus, we believe these two factors contribute to the situation that systematic research on standalone 16-bit training remains notably absent, a critical gap for hardware-constrained ML practitioners.
In this research, we aim to fill in this understudied area, standing from the perspective of resource-constrained ML practitioners to analyze standalone 16-bit training—both theoretically and practically. Here, we define "standalone" as employing 16-bit precision exclusively for training, weight storage, and inference, without the support of higher or lower precision formats in any phase. For benchmarking, we will compare standalone 16-bit training against 32-bit training and mixed precision methods. Our objective is to assess whether 16-bit training can independently serve as a feasible option. We believe that the results of this investigation could substantially improve the accessibility and utility of machine learning across a wider array of individuals and organizations.
Figure 1: Comparison of 16-bit floating-point neural networks with other precisions across different image classification architectures on CIFAR-10. Circle sizes indicate the number of parameters in each architecture. 16-bit, mixed precision, and 32-bit are represented by red, yellow, and blue colors, respectively. The x-axis denotes time, and the y-axis denotes accuracy.
@@
Our work proceeds as follows. We start with a theoretical examination of the differences in accuracy between 16-bit and 32-bit. This theory includes formalizing concepts related to error tolerance and floating-point errors, proposing and proving a lemma that theoretically ensures that 16-bit can achieve comparable classification results to 32-bit under specific conditions. Preliminary observations support the hypothesis that 16-bit and 32-bit should perform similarly in handling classification problems. Then, through numerous comparisons against 32-bit and mixed-precision training, we highlight the relevance of standalone 16-bit training in image classification with CNN and ViT [13] architectures.
Fig. 1 illustrates some of our experimental results. All results were obtained using SGD optimizers rather than the default Adam optimizer, which likely contributed to the perception of 16-bit precision as insufficient for standalone use and, consequently, less studied (as mentioned above). To our knowledge, these results represent the first comprehensive evidence to suggest that standalone 16-bit floating-point neural networks can perform on par with their mixed-precision and 32-bit counterparts in classification. These findings should align with the expectations of many ML practitioners. However, it is essential to back such claims with scientific evidence to further the field. This work provides that missing piece of evidence through both theoretical and empirical perspectives.
In summary, this paper offers the following contributions:
  • Theoretical insights into the accuracy differences between 16-bit and 32-bit. Our theory provides a viable alternative to the perception noise in neural networks may act as regularization.
  • An extensive experimental study comparing 16-bit with 32-bit and mixed precision. We present a detailed empirical investigation to demonstrate the extent to which 16-bit floating-point can function as a standalone.
  • Identification of factors influencing 16-bit models. This paper further examines the factors that could influence the effective use of 16-bit models.
To facilitate reproducibility, we provide detailed descriptions of our experimental settings in the paper. We will release our code and evaluation results to the public upon publication.
@@ 2 Related Work
In the past decade, we have observed that scaling networks significantly improves their abilities to resolve complex tasks [31]. These increasingly large networks require extensive computations for their training and large-scale deployment. This trend raises several issues regarding accessibility, reproducibility, efficiency, environmental impact, and real-time execution. Given these critical challenges, numerous approaches have been proposed to address them and offer solutions to reduce energy and memory footprints while decreasing both running and training time. Quantization techniques reduce the precision of weights and activations to improve efficiency, as demonstrated by Jacob et al. [26], while methods like PACT [4] further optimize performance by learning clipping ranges. Pruning, as explored by Han et al. [20], removes redundant parameters to compress models, and Gale et al. [15] have shown its effectiveness in reducing resource usage without significant accuracy loss. Distillation, introduced by Hinton et al. [24], transfers knowledge from large models to smaller ones, and Touvron et al. [52] have adapted it for transformers. While these techniques are often viewed as primary solutions to these challenges, they are typically applied as post-processing steps before model deployment. Consequently, they do not reduce the training time and necessitate an additional step before deployment. In this work, we focus on the specific case of standalone IEEE 16-bit floating-point neural networks for both training and inference without any additional stages. Thus, this literature review will exclusively concentrate on works designed to accelerate both of these stages.
Low precision training pipeline. Training neural networks with low-bit precision while maintaining accuracy has been a critical focus in the literature. A key approach involves adapting the training pipeline to efficiently work with low-precision formats. For instance, early work [11] showed that using low-precision multiplications (e.g., 10-bit) can yield results comparable to 32-bit training, thanks to selective precision adjustments for different stages of computation. However, this approach can introduce challenges related to numerical stability and the risk of gradient underflow during training. Recent work such as LLM.int8() [12] introduces efficient 8-bit matrix multiplication techniques for transformers, significantly reducing computational cost while preserving accuracy. Nonetheless, this method requires careful handling of sensitive layers in the model, such as attention mechanisms, where quantization errors may arise.
At the core of low-precision approaches are two main types of numerical formats: floating-point and fixed-point. Floating-point numbers, commonly used in deep learning, can represent a wide range of values, which is important for complex tasks that require high precision. Fixed-point formats, on the other hand, use a fixed allocation of bits for integer and fractional parts, offering faster computation and lower memory usage, but at the cost of reduced precision [43]. This makes fixed-point formats more suitable for low-power, resource-constrained systems such as embedded devices and edge computing, although they may struggle with high-precision tasks that demand a wider dynamic range.
In the context of low-precision training, fixed-point strategies modify data representations while keeping the overall training pipeline unchanged. This simple and efficient approach is ideal for systems with limited computational resources. However, fixed-point formats often face challenges with range and precision, which can hinder their effectiveness in tasks that require finer numerical accuracy. Studies like [19], [3], and [36] have explored how fixed-point precision can lead to accuracy drops in neural networks and have proposed refinements to minimize these losses while maintaining efficiency. Nevertheless, these solutions may not always fully resolve the balance between efficiency and precision, especially in larger, more complex models where accuracy is more sensitive to the precision of weights and activations.
Moreover, integer quantization, a variant of fixed-point representation, has become an effective method for reducing computation costs. Several works, including [54], [10], and [26], have demonstrated the utility of integer quantization for neural network weights and activations. Despite its advantages in reducing computational overhead, integer quantization can lead to degradation in model accuracy, particularly in more complex models where maintaining sufficient precision is crucial. Recent advancements like SmoothQuant [56] and FlexPoint [33] dynamically adapt precision formats to further optimize memory and performance during training, but these approaches also require careful tuning to avoid accuracy loss. For example, studies such as Q8BERT [57] have applied 8-bit quantization to transformer models like BERT, demonstrating significant reductions in computational complexity. However, TernaryBERT [58], which applies 3-bit quantization to transformer models, reduces model size and energy consumption but presents challenges related to maintaining model accuracy, particularly in more detailed tasks, due to the lower bit-width.
Our research proposes a simple method that utilizes IEEE 16-bit floating-point operations for both training and inference of neural networks, without the need for complex integer quantization or dynamic application of various formats.
Mixed precision. The inception of a practical low-precision floating-point mechanism can be traced back to the work of Courbariaux et al. [8], who devised BinaryConnect, a method that reduces memory usage by binarizing weights during training while maintaining 32-bit precision for gradient updates, making it suitable for simpler models but potentially less effective for more complex ones. XNOR-Net [46] extended this by introducing binary operations into CNNs, further reducing computation but potentially sacrificing some accuracy in more detailed tasks, especially in vision tasks. Similarly, Jacob et al. [27] proposed an integer-only quantization method that applies quantization in the forward pass while retaining floating-point precision during backpropagation, balancing computational efficiency and model accuracy. However, the use of lower bit-widths may require careful calibration to avoid accuracy loss in more complex models.
Zhou et al. [59] introduced DoReFa-Net, which employs extremely low-bit representations (e.g., 2-bit activations and 4-bit weights) to minimize power consumption and memory usage, making it ideal for edge devices. However, the lower bit-widths require additional fine-tuning to maintain acceptable performance in complex tasks, and extreme quantization may result in significant accuracy loss if not calibrated correctly.
Micikevicius et al. [38] proposed a mixed precision training approach using 16-bit forward and backward propagations while maintaining 32-bit weights for updates. While this reduces memory usage, additional storage and conversion processes can sometimes offset the performance improvements. Wang et al. [53] proposed an 8-bit training method combined with 16-bit accumulations, significantly reducing computational load while preserving accuracy. Yet, as model complexity increases, ensuring sufficient precision can become challenging. Lastly, Koster et al. [29] revisited BF16, a 16-bit floating-point format with a broader dynamic range compared to FP16, making it more effective for deep learning tasks without significant accuracy loss. BF16 has gained widespread use in platforms like TPUs due to its ability to balance precision and performance in large-scale models.
Our approach differs by focusing solely on the use of IEEE 16-bit precision for both training and inference, without the need for mixed-precision adjustments or additional complexity, demonstrating that neural networks can achieve competitive accuracy using 16-bit precision alone.
Hardware support. Sharma et al. [50] introduced BitFusion, a dynamic hardware accelerator that adjusts bit-level operations at runtime to optimize resource use. This adaptive precision across neural network layers leads to significant energy savings and efficiency, making it ideal for low-precision tasks, though managing varying precision levels can introduce hardware design complexity. Similarly, fixed-point arithmetic, which reduces precision to cut computation and memory needs, is supported by hardware like FPGAs and ASICs [43]. In resource-constrained environments, such as embedded systems, fixed-point DSP processors are commonly used to accelerate low-precision operations [47].
Google’s Tensor Processing Unit (TPU) supports BF16, improving performance and memory efficiency in large-scale training [29]. Recently, NVIDIA’s Ampere and Hopper architectures expanded support for FP16, BF16, FP8, and FP4, further enhancing low-precision efficiency for both deep learning and large-scale models such as transformers [39]. While general-purpose GPUs support low-precision formats, specialized hardware such as BitFusion, TPUs, and fixed-point devices (e.g., FPGAs, DSPs) are essential for maximizing the benefits of low-precision computing, especially in large-scale training and edge applications. These technologies often require specific hardware environments optimized for energy efficiency and computational throughput.
@@ 3.3 Observation: Standalone IEEE 16-bit Floating-Point DNN on MNIST
We validate Lemma 3.3 by comparing the performance of a standalone IEEE 16-bit floating-point deep neural network to a 32-bit one on MNIST. Our implementation is done using TensorFlow. The 16-bit implementation follows the same architecture as the 32-bit counterpart, except that all floating-point operations are performed using 16-bit precision.
@@
This finding motivates us to investigate whether the similarity between the IEEE 16-bit and 32-bit models holds true for more complex neural networks. In theory, if a certain percentage, denoted by $a\%$, of samples in a dataset satisfy Eq. 3, Lemma 3.3 guarantees that the 32-bit and 16-bit models will have at least $a\%$ of their classification results in common. However, determining the exact $a\%$ poses a challenge. Nevertheless, we cautiously believe that for complex neural networks, a large majority of data satisfies Eq. 3. This belief stems from the fact that the loss function for classification problems, in the form of cross-entropy $-\Sigma\log(p_{i})$, guides the probabilities $p_{i}$ towards 1 during training, thereby resulting in a larger error tolerance $\Gamma$ compared to the relatively smaller $\delta$.
@@ 4 Experiments
Result Summary
IEEE-16bit floating-point demonstrates standalone potential.
  • Ease of implementation: IEEE-16bit can be straightforwardly integrated, replacing 32-bit operations with 16-bit ones (Sect. 4.1). Generally, no fine-tuning is necessary (Sect. 5).
  • Accuracy: IEEE-16bit outperforms 32-bit in 3 out of the 12 examined architectures and exceeds Mixed precision in a 3 of them. On average, the accuracy of IEEE-16bit is slightly less than the accuracy of 32-bit and Mixed precision, 0.2% (Tab. 2). An everyday machine learning user may not notice distinction in accuracy (Fig. 4).
  • Runtime Efficiency: IEEE-16bit processes on average 1.6 times faster than 32bit and has a 1.3 times speed advantage over Mixed precision (Tab. 2).
@@
4.1 Baseline Precisions
We define three precision schemes denoted by FP16, FP32, and MP, respectively, which form the basis of our comparative study.
FP16:
This refers to the IEEE 16-bit floating-point precision. Our implementation is achieved from TensorFlow using tf.keras.backend.set_floatx(’float16’)[18]. Because operations involving both 16-bit and 32-bit can implicitly cast 16-bit values to 32-bit, we have cautiously examined the data types of the weights at various before, during, and post-training to ascertain they are indeed 16-bit and avoid unintentional mixture with 32-bit.
FP32:
This represents the conventional 32-bit floating point precision as defined by the IEEE standard. We utilize TensorFlow’s default implementation for this precision level.
MP:
The Mixed Precision scheme we refer to here is consistent with TensorFlow’s approach [18]. Our implementation is achieved from TensorFlow using tf.keras.mixed_precision.set_global_policy(’mixed_float16’). This methodology, initially introduced by Baidu and NVIDIA [38], is also supported in PyTorch, as highlighted in NVIDIA’s 2023 report [40].
4.2 Datasets and Architectures
For our image classification experiments, we employ the CIFAR-10 dataset [30], which consists of 60000 32x32 colour images in 10 classes. By default, this dataset is in 32-bit, but we convert it to IEEE 16-bit precision before training with FP16. Our experiments encompass 12 architectures: 10 established CNN architectures and the state-of-the-art Vision Transformer (ViT-8 and ViT-12) [13]. These CNN models include AlexNet [32], VGG16, VGG19 [51], ResNet-34, ResNet-56 [21], MobileNetV2 [49], DenseNet-121, DenseNet-169 [25], and Xception [5]. We don’t use pre-train weights or models for our experiments.
4.3 Measurement Metrics, Software and Hardware Settings
We carried out our experiments using 50 different random seeds, averaging both accuracy and training times to guarantee unbiased results. We carried out our experiments using 50 different random seeds, averaging both accuracy and training times(wall-clock-time) to guarantee unbiased results. Our tests spanned 100 epochs each, maintaining a consistent batch size of 256 and learning rate of 0.01 with the SGD optimizer [17] and momentum 0.9 [45]. We opted for the default settings, avoiding any parameter adjustments, and excluded data augmentation from our processes. All experiments have been conducted on a RTX4080 GPU and 40GB of RAM.

### v6

# Corresponding author. †Equal contribution

@@ Abstract
Keywords:
Neural Networks Deep Learning Low Precision Efficient Deep Learning
@@ 1 Introduction
As machine learning models grow in complexity and size, the demand for computational resources such as memory and processing power has become a critical challenge. To address this, the mixed precision technique has been developed to balance resource consumption and computational efficiency [33, 48, 6, 50]. This approach strategically employs different numerical precision types during forward and backward propagation, weight storage, or inference. For instance, GPT-3 was trained with mixed precision of 32-bit and 16-bit formats [2], while Meta’s LLaMA 3 employs FP8 precision in both inference and forward pass training phases [12, 39].
The trend toward mixed precision is closely linked to the development of lower-precision numerical formats. The ongoing ML arithmetic standard IEEE P3109 has formalized seven 8-bit floating-point formats [1], and the Nvidia Blackwell chip is pioneering 4-bit precision [32, 30]. However, while mixed precision is designed to mitigate resource pressures, access to this technology often remains elusive for many machine learning practitioners facing hardware limitations.
Table 1 shows major Nvidia GPUs and their hardware support for floating-point operations. Both FP32 and FP16 are supported across all GPUs listed, while support for newer formats like bfloat16 and FP8 is limited to recent architectures [36, 37]. For practitioners constrained by hardware, the options for training models are typically limited to using 32-bit, 16-bit, or their combination in mixed precision.
Table 1: Supported floating-point precisions across major NVIDIA GPU architectures, showing both CUDA Core and Tensor Core capabilities. [36, 37]
@@
Among these options, 16-bit precision should theoretically be preferred, as it offers memory and performance advantages while being well-supported by modern hardware. Yet, surprisingly, there’s limited information about the ability of 16-bit formats to work independently. This work focuses on IEEE binary16 due to its wider accessibility compared to Google Brain’s bfloat16.
The absence of research on standalone 16-bit functionality can be attributed to two factors: the shift towards even lower precision models by those with advanced hardware, and the observation of numerical instability in early experiments that led to mixed-precision solutions [48]. While direct attempts to train networks using 16-bit precision often fail in frameworks like TensorFlow or PyTorch, we show these issues stem from incorrect parameter settings, particularly in the Adam optimizer’s epsilon value.
@@
Our work proceeds as follows. We start with a theoretical examination of accuracy differences between 16-bit and 32-bit formats, proposing a lemma that ensures comparable classification results under specific conditions. Through extensive comparisons against 32-bit and mixed-precision training, we demonstrate the viability of standalone 16-bit training in image classification using CNN and ViT [11] architectures. To ensure reproducibility, we provide detailed experimental settings and will release our code upon publication.
@@ 2 Related Work
In the past decade, we have observed that scaling networks significantly improves their abilities to resolve complex tasks [27]. These increasingly large networks require extensive computations for their training and large-scale deployment. This trend raises several issues regarding accessibility, reproducibility, efficiency, environmental impact, and real-time execution. Given these critical challenges, numerous approaches have been proposed to address them and offer solutions to reduce energy and memory footprints while decreasing both running and training time. Quantization techniques reduce the precision of weights and activations to improve efficiency, as demonstrated by Jacob et al. [23], while methods like PACT [4] further optimize performance by learning clipping ranges. Pruning, as explored by Han et al. [17], removes redundant parameters to compress models, and Gale et al. [13] have shown its effectiveness in reducing resource usage without significant accuracy loss. Distillation, introduced by Hinton et al. [21], transfers knowledge from large models to smaller ones, and Touvron et al. [47] have adapted it for transformers.
Low Precision Training Pipeline. Training neural networks with low-bit precision while maintaining accuracy has been a critical focus in the literature. Early work [9] showed that using low-precision multiplications can yield results comparable to 32-bit training, thanks to selective precision adjustments. Recent work such as LLM.int8() [10] introduces efficient 8-bit matrix multiplication techniques for transformers, though requiring careful handling of sensitive layers where quantization errors may arise. At the core of low-precision approaches are two main types: floating-point and fixed-point formats. Floating-point numbers, commonly used in deep learning, can represent a wide range of values crucial for complex tasks requiring high precision. In contrast, fixed-point formats [38] use a fixed allocation of bits for integer and fractional parts. While offering faster computation and lower memory usage, fixed-point formats face fundamental challenges with range and precision, particularly struggling with tasks that demand wider dynamic range. Studies like [16], [3], and [31] demonstrate how these limitations lead to accuracy drops in neural networks, especially in larger, more complex models where accuracy is highly sensitive to weight and activation precision. Integer quantization approaches, including [49], [8], and [23], attempt to address these challenges but still face accuracy degradation in complex models. Even recent advances like SmoothQuant [51] and FlexPoint [29] require careful tuning. For example, while Q8BERT [52] shows promise with 8-bit quantization, TernaryBERT [53] faces significant accuracy challenges due to its lower bit-width. Our research proposes a simple yet effective method using IEEE 16-bit floating-point operations for both training and inference, avoiding the inherent limitations of fixed-point representations while maintaining the precision necessary for modern deep learning.
Mixed Precision. The inception of practical low-precision mechanisms traces back to Courbariaux et al. [7], who devised BinaryConnect. XNOR-Net [41] extended this by introducing binary operations into CNNs. Zhou et al. [54] introduced DoReFa-Net for edge devices. Micikevicius et al. [33] proposed mixed precision training using 16-bit operations while maintaining 32-bit weights. Wang et al. [48] developed an 8-bit training method with 16-bit accumulations. Koster et al. [25] revisited BF16, making it effective for deep learning tasks without significant accuracy loss.
Hardware Support. Sharma et al. [44] introduced BitFusion, a dynamic hardware accelerator optimizing resource use. Fixed-point arithmetic is supported by FPGAs and ASICs [38], and fixed-point DSP processors are commonly used in embedded systems [42]. Google’s TPU supports BF16 [25], while NVIDIA’s recent architectures expanded support for various precisions [34]. While general-purpose GPUs support low-precision formats, specialized hardware such as BitFusion, TPUs, and fixed-point devices are essential for maximizing the benefits of low-precision computing.
@@ 3.3 Observation: Standalone IEEE 16-bit Floating-Point DNN on MNIST
We validate Lemma 1 by comparing the performance of a standalone IEEE 16-bit floating-point deep neural network to a 32-bit one on MNIST. Our implementation is done using TensorFlow. The 16-bit implementation follows the same architecture as the 32-bit counterpart, except that all floating-point operations are performed using 16-bit precision.
@@
This finding motivates us to investigate whether the similarity between the IEEE 16-bit and 32-bit models holds true for more complex neural networks. In theory, if a certain percentage, denoted by $a\%$, of samples in a dataset satisfy Eq. 3, Lemma 1 guarantees that the 32-bit and 16-bit models will have at least $a\%$ of their classification results in common. However, determining the exact $a\%$ poses a challenge. Nevertheless, we cautiously believe that for complex neural networks, a large majority of data satisfies Eq. 3. This belief stems from the fact that the loss function for classification problems, in the form of cross-entropy $-\Sigma\log(p_{i})$, guides the probabilities $p_{i}$ towards 1 during training, thereby resulting in a larger error tolerance $\Gamma$ compared to the relatively smaller $\delta$.
@@ 4 Experimental Results
4.1 Precisions
We compare three floating-point precision schemes: FP16 (IEEE 16-bit), FP32 (standard 32-bit), and MP (mixed precision). For FP16, we use TensorFlow’s tf.keras.backend.set_floatx(’float16’) and ensure no unintended type casting to FP32 occurs during training. FP32 uses TensorFlow’s default 32-bit precision. For MP, we follow TensorFlow’s mixed precision policy via tf.keras. mixed_precision.set_global_policy(’mixed_float16’), which allows FP16 operations with FP32 accumulations [33, 15, 35], as originally proposed by Baidu and NVIDIA.
4.2 Experimental Settings
We conduct image classification experiments using the CIFAR-10 dataset [26], converting it from 32-bit to IEEE 16-bit for FP16 training. Our evaluation spans 12 architectures, including 10 popular CNNs (AlexNet [28], VGG16/19 [45], ResNet-34/56 [18], MobileNetV2 [43], DenseNet-121/169 [22], and Xception [5]) and two Vision Transformers (ViT-8 and ViT-12 [11]). All models are trained from scratch without pre-trained weights. Each experiment runs for 100 epochs with a batch size of 256 and a learning rate of 0.01 using SGD [14] with momentum 0.9 [40]. No data augmentation or hyperparameter tuning is applied. Results are averaged over 50 random seeds, and all experiments are performed on a single RTX 4080 GPU with 40GB RAM, reporting both accuracy and wall-clock training time.